OpenStudy (anonymous):

a car is moving at a velocity of 20m\s. how long will it take to travel 1km

OpenStudy (anonymous):

\[v = t/l\]

OpenStudy (anonymous):

convert m/s to km/h, and then figure it out from there :)

OpenStudy (shane_b):

1km = 1000m so you can just do this:\[\frac{1000m}{20m/s}=50s\]